### v1

# Uncertainty Propagation, and Parameter Calibration

Abstract
The goal of this paper is to demonstrate and address challenges related to all aspects of performing a complete uncertainty quantification (UQ) analysis of a complicated physics-based simulation like a 2D slab burner direct numerical simulation (DNS). The UQ framework includes the development of data-driven surrogate models, propagation of parametric uncertainties to the fuel regression rate–the primary quantity of interest–and Bayesian calibration of critical parameters influencing the regression rate using experimental data. Specifically, the parameters calibrated include the latent heat of sublimation and a chemical reaction temperature exponent. Two surrogate models, a Gaussian Process (GP) and a Hierarchical Multiscale Surrogate (HMS) were constructed using an ensemble of 64 simulations generated via Latin Hypercube sampling. Both models exhibited comparable performance during cross-validation. However, the HMS was more stable due to its ability to handle multiscale effects, in contrast with the GP which was very sensitive to kernel choice. Analysis revealed that the surrogates do not accurately predict all spatial locations of the slab burner as-is. Subsequent Bayesian calibration of the physical parameters against experimental observations resulted in regression rate predictions that closer align with experimental observation in specific regions. This study highlights the importance of surrogate model selection and parameter calibration in quantifying uncertainty in predictions of fuel regression rates in complex combustion systems.
1 Introduction
Hybrid rocket systems, characterized by a fuel and oxidizer in two different states, are considered better alternatives compared to other propulsion systems due to their ability to operate with a highly dense fuel source similar to a bi-propellant solid motor but with the operational advantages of a liquid motor [1, 2, 3]. Many experimental studies have been conducted that test slab burner hybrid rocket setups with high alkane fuels such as paraffin because these fuels achieve higher regression rates compared to other options (e.g., see [4, 5, 6, 7, 8, 9]). The higher regression rates result from the combustion phenomena of such fuels: the formation of a liquid layer on the surface of the solid, which, together with the formation of instabilities in the fuel-oxidizer interface, leads to the entrainment of combustible liquid droplets into the main flow [10]. Despite the observation of these concepts experimentally, creating accurate computational models for hybrid rockets is very challenging due to the complexity of the physics involved: flow transport, atomization, turbulent combustion, and complex chemical kinetics need to be solved together. Accurate prediction of reacting flows, in general, is a challenge in predictive modeling because of the multiscale nature of combustion, as the underlying coupled physical phenomena occur under a wide range of time and length scales. Namely, elementary chemical reaction processes happen in short time scales (e.g., chain reactions that occur on the order of $10^{-10}s$), whereas the scales of the flow transport are much larger on the order of $10^{-4}$ to $10^{-2}s$  [11]. Similarly, in terms of length dimension scales, the flame structures under radiation and soot effects are smaller than physical scales of flow transport in a combustion chamber  [12, 13]. Apart from combustion being multiscale, there exists varying degrees of coupling between the physical processes involved. For example, chemical reactions and molecular diffusion are tightly coupled at fine scales [14] and need to be modeled e.g., via reduced-order flamelet manifolds (see [15, 16, 17, 18]) or stochastically with probability distributions [19, 20] in turbulent combustion simulations since grids chosen to capture flows are not sufficiently fine.
These challenges of the coupled multiscale phenomena have led to the development of various computational approaches, such as Direct Numerical Simulation (DNS). In DNS, the Navier-Stokes equations are coupled with the thermochemical species equations and solved numerically on a grid that sufficiently captures the underlying scales [21]. An issue that arises when using DNS solvers is that the computational cost is often prohibitive for any detailed data-driven analysis that requires many simulation runs like uncertainty quantification (UQ). DNS simulations of combustion are often not standalone components, but parts of a larger framework that may include experiments and/or other models that either provide inputs to the DNS or use the DNS outputs. In this context, a careful UQ analysis on the DNS is required to either validate the DNS solvers themselves for decision-making depending on the application or to be able to propagate uncertainty through the DNS to other models. Any standard forward UQ or calibration analysis requires large ensembles of model evaluations to adequately approximate the true posterior distribution of the output quantities of interest (QoIs) in forward UQ or the posterior of the calibration parameters [22, 23], which is infeasible given the DNS computational cost. Therefore, it is common practice and a requirement to create low-cost and high-accuracy surrogate models (i.e., emulators, approximate models) [24, 25, 26] for uncertainty quantification and calibration of these expensive solvers.
There are many methods for developing surrogate models. Polynomial chaos expansions (PCE) approximate the input to QoI function $\mathcal{M}:\bm{x}\in\mathcal{D}_{\bm{X}}\subset\mathbb{R}^{M}\mapsto\bm{y}=\mathcal{M}(\bm{x})$ with polynomial approximations [27]. Canonical low-rank tensor approximations (LRA) approximate the input to QoI function with products of univariate functions that can also be orthogonal polynomials similar to PCE [28]. Feedforward neural networks map inputs to QoIs via learning the weights and biases between layers of neurons [29, 30]. Kriging or Gaussian Processes (GPs) assume the likelihood function of the QoI to be a multivariate normal distribution modeled as the addition of a mean regression term and a Gaussian stationary random process [31]. Combinations of these methods are also options for surrogate models, e.g., Deep GPs [32], where the data between layers of a neural network are modeled as a multivariate GP. Traditional surrogate methods such as the ones we outline above often perform poorly when used in multiscale problems, because these problems include characteristics that are computationally challenging for these basic methods. Firstly, multiscale problems have a very large input space and that requires specific techniques to address the curse of dimensionality before proceeding with surrogate models [33, 34, 35]. Secondly, methods such as GPs are not scalable because they include covariance matrix inversions, which are computationally expensive. Thirdly, the assumptions of smoothness and stationarity imposed by Gaussian assumptions or polynomials does not necessarily hold for multiscale problems. Lastly, the ability of simpler methods to capture non-linear trends from high-dimensional data is limited [36].
The main goal of this paper is to systematically address challenges related to all aspects of performing a complete UQ analysis of a complicated physics-based simulation like a 2D slab burner DNS. We provide insights related to the development of an acceptable surrogate model, the propagation of uncertainty of the DNS inputs to the QoI, and the calibration of DNS parameters based on observations of experimental data. The DNS used in this work is ABLATE (Ablative Boundary Layers At The Exascale), an in-house DNS solver developed for simulating a multiphase 2D slab burner setup to better understand combustion in hybrid rocket motors, which includes ablation, radiation, and soot modeling[37, 21, 15]. ABLATE uses the PETSc libraries from Argonne National Laboratory [38, 39] for data management, and parallel large scale computation. For the work described in this paper, ABLATE simulated the 2D reacting flow between pure oxygen $O_{2}$ (the oxidizer) and Methyl methacrylate MMA (the fuel) with chemistry modeled by a reduced mechanism. To accurately capture the multiscale behavior and overcome the computationally prohibitive UQ analysis of the DNS simulations, we build both GPs and a novel computationally efficient hierarchical multiscale surrogate (HMS) [40]. HMS is based on a forward-backward greedy approach to construct a set of data driven basis functions with a multiscale structure. Forward-backward here means the iterative selection of basis functions at different scales. This approach generates hierarchical basis functions belonging to Reproducing Kernel Hilbert Spaces (RKHS), and the hierarchical models are able to accurately represent irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting by using kernels at successively finer scales. We conduct a forward UQ analysis using the developed surrogate and uncertainty in a subset of inputs identified using sensitivity analysis and propagating the uncertainty of the DNS inputs to the slab burner output QoI. The inputs to the DNS are the velocity factor for the far-field inlet oxidizer flux assuming a parabolic profile ($V_{in}$), the thermochemical properties of the chemistry mechanism ($\theta_{c}$: activation energies $E_{*}$, temperature exponents $b_{*}$, and pre-exponential factors $A_{*}$), the fixed geometry of the slab burner, and the fuel parameters ($\theta_{f}$). Combustion DNS solvers produce a variety of outputs that may be of interest depending on the goals. Here, we are most interested in QoIs along the boundary of the slab, namely the regression rate ($\dot{r}$). The regression rate has a direct impact on the geometrical design of the rocket motor and its performance ([41], [42]). Lastly, we leverage the constructed surrogate model and experimental regression rate measurements to calibrate the fuel ($\theta_{f}$) and chemistry ($\theta_{c}$) parameters of the DNS. In particular, we use the Bayesian method to infer the probability distribution of the latent heat of sublimation ($l_{v}$) and temperature exponent for a particular important reaction ($b_{249}$).
The paper is organized as follows. Section 1 acts as the introduction and provides motivation for uncertainty quantification of hybrid rocket DNS solvers. Section 2 presents ABLATE, the DNS solver for a 2D slab burner with descriptions of its inputs and outputs. Section 3 introduces the hierarchical multiscale surrogate (HMS) and benchmarks, as well as the forward UQ and calibration methods. Section 4 includes results and discussions for the development of the surrogate models, forward UQ, and calibration. Section 5 is the concluding section with directions for future work.
2 ABLATE: Coupled Flow/Combustion DNS
In this section, we describe our coupled flow/combustion solver ABLATE (https://ablate.dev/), its inputs and outputs, and how we used it to generate the dataset to train the surrogates in Section 3. The solver simulates a 2D slab burner setup for the combustion of MMA as the fuel and pure oxygen $O_{2}$ as the oxidizer. Figure  1 shows the two-dimensional slab burner domain. An inlet of pure oxygen flows from the left side over the fuel slab where it reacts with the fuel vapors and flows through the outlet of the slab burner on the right side. There are four different boundary conditions used in the simulation depending if the boundary is a (1) wall, (2) inlet, (3) outlet, or (4) the fuel. The gas phase system for the slab burner environment is described by the reacting compressible Navier-Stokes equations, coupled with radiation heat transfer:
$\frac{\partial\rho}{\partial t}+\frac{(\partial\rho u_{i})}{\partial x_{i}}=0$
$\frac{\partial\rho u_{j}}{\partial t}+\frac{\partial(\rho u_{i}u_{j})}{\partial x_{i}}=-\frac{\partial p}{\partial x_{j}}+\frac{\partial\tau_{ij}}{\partial x_{i}}$
$\frac{\partial(\rho Y_{k})}{\partial t}+\frac{\partial(\rho u_{i}Y_{k})}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(\rho D_{k}\frac{\partial Y_{k}}{\partial x_{i}}\right)+\dot{m}^{\prime\prime\prime}_{k}$ (1)
$\frac{\partial(\rho e_{t})}{\partial t}+\frac{\partial(\rho u_{i}H_{t})}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(u_{i}\mu\frac{\partial u_{i}}{\partial x_{i}}\right)+\frac{\partial}{\partial x_{i}}\left[\left(k\frac{\partial T}{\partial x_{i}}+\sum^{N}\rho D_{k}\frac{\partial Y_{k}}{\partial x_{i}}h_{k}\right)\right]$
$\hskip 91.04872pt-\sum\dot{m}^{\prime\prime\prime}_{k}h_{f,k}^{\circ}-\frac{\partial}{\partial x_{i}}(\rho Y_{C(s)}V_{T,i}h_{C(s)})-\kappa_{P}(\sigma T^{4}-G)$
where, $k$ is the thermal conductivity, $\dot{m}^{\prime\prime\prime}_{k}$ is the mass consumption or production rate of the $k^{th}$ species from chemical reactions, $e_{t}=e+u_{i}^{2}/2$ is the total sensible energy, $H_{t}=e_{t}+p/\rho$ is the total enthalpy and $h_{f,k}^{\circ}$ is the heat of formation of the $k^{th}$ species. $V_{t,x}$ is the thermophoretic velocity associated with soot thermophoresis. A SutherLand transport model is used along with a unity Lewis number assumption. The species diffusivities are assumed to be the same for all species. The chemical kinetic mechanism for MMA oxidation, developed by Bolshova et. al. [43], is utilized consisting of $67$ species and $263$ reactions. A six step semi-empirical soot model based on an acetylene precursor is adapted that includes rates for nucleation, agglomeration, surface growth and oxidation [44].
Figure 1: A sample temperature contour of the computational domain showing the specified inlet, outlet, isothermal wall, and fuel boundaries. The QoI is the time averaged regression rate at each point along the fuel boundary.
[…]
The wall boundary conditions are simply treated as isothermal no-slip walls held at the ambient temperature of 300 K. The inlet and outlet boundary conditions are described using locally one-dimensional inviscid characteristics (LODI) [48]. The velocity profile prescripted at the inlet is described using a $1/7^{th}$ power law for fully developed turbulent profile for pipe flow, $u_{in}(r)=V_{in}(1-2r/D_{pipe})^{1/7}$, where $V_{in}$ is a velocity factor defined by the oxidizer mass flux to obtain the correct mass flow rate, $A=60~{}G/49\rho_{ox}$.
The fuel boundary energy balance is used to couple the heat flux incident on the fuel surface from the gas phase to the mass of fuel vaporizing of the surface,
$\dot{q}^{\prime\prime}_{c,g}+\dot{q}^{\prime\prime}_{r,Net}=\dot{m}^{\prime\prime}_{f}L_{v}+\dot{q}^{\prime\prime}_{c,f}.$ (2)
where $\dot{q}^{\prime\prime}_{c,g}(=k\vec{\nabla}T\cdot\hat{n}_{s})$ is the conductive heat flux from the gas phase normal to the fuel surface, $\dot{q}^{\prime\prime}_{c,f}$ is the conductive heat flux used in heating the fuel, and $\dot{q}^{\prime\prime}_{r,Net}(=\alpha_{s}\dot{q}^{\prime\prime}_{r,abs}-\epsilon_{s}\sigma T_{f}^{4}$) is the net radiation heat transfer at the fuel surface. $\dot{q}^{\prime\prime}_{r,abs}$ is the irradiant radiation heat flux onto the fuel surface, and $T_{f}$ is the fuel surface temperature. $\epsilon_{s}$ and $\alpha_{s}$ are the fuel surface emissivity and absorptivity. $\dot{m}^{\prime\prime}_{f}=\rho_{f}\dot{r}$ is the local fuel mass flux and is related to the local regression rate by the solid fuel density $\rho_{f}$. The current simulations assume the fuel is isothermal and already at the vaporization temperature of PMMA ($T_{f}=653~{}K$) such that the is no heat loss due to heating the fuel ($\dot{q}_{c,f}=0$). This leads to the following mass flux and regression rate relations depending only on the instantaneous gas phase solution,
$\dot{r}(x,y,t)=\frac{\dot{m}^{\prime\prime}_{f}(x,y,t)}{\rho_{f}}=\frac{k\vec{\nabla}T|_{s}(x,y,t)\cdot\hat{n}_{s}+\alpha_{s}\dot{q}^{\prime\prime}_{r,abs}(x,y,t)-\epsilon_{s}\sigma T_{f}^{4}}{\rho_{f}L_{v}}$ (3)
In this isothermal fuel boundary model, the fuel temperature is held at $T_{f}=653K$. The density, MMA species mass, and normal momentum boundary conditions have additional mass and momentum fluxes due to $\dot{m}^{\prime\prime}_{f}$.
The system is solved using the DNS framework (ABLATE) which is available under a BSD-3-Clause License hosted on GitHub.com/UBCHREST/ablate. ABLATE utilizes a finite volume formulation on an unstructured grid employing the AUSM family of flux vector splitting schemes to solve the presented conservation equations. Time advancement is achieved using high-order Runge-Kutta methods. Multi-block unstructured mesh domain decomposition is employed for high-efficiency parallel computation built upon PETSc’s DMPlex [49]. Additional libraries include the open source library TChem [50], which is used to integrate the chemical kinetics required for the combustion modeling.
The QoI the surrogates are trained on is a time-step weighted average of the regression rate at each location on the boundary $\mathcal{S}_{B}$. The QoI is selected as such to be able to later use the equivalent measurement from experimental data (regression rate between image snapshots) for calibration.
[…]
Where $\mathcal{S}_{B}$ includes the paired coordinates $x,y$ for the grid points on the boundary, $N$ is the number of simulation time steps, $t_{i}$ is the elapsed simulation time, and $\dot{r}(x,y,t)$ is the regression rate as computed during the simulation from Eqn. 3. Each simulation of the DNS resolves, on average, about 25-75ms of the reacting flow (1 to 3 flow through times) and takes 24 hours to complete.
3 Surrogate Models and Uncertainty Quantification Methods
This section provides details of the surrogate models (Sections 3.1–3.2), the sampling ensemble strategy to train them (Section 3.3), and their implementation for facilitating forward UQ (Section 3.4) and Bayesian calibration (Section 3.5). All the data come from the DNS 2D slab burner reacting flow simulation described in the previous section.
3.1 Gaussian Processes (GPs)
GPs are widely used surrogates because they are fast, and therefore allow for UQ analyses of expensive computational models to be feasible. GP assumes that the likelihood function of the QoI is a multivariate normal distribution (Eq. 5).
$[y_{x,y}(x_{1}),y_{x,y}(x_{2}),...,y_{x,y}(x_{n})]^{T}\approx\mathcal{MN}([\mu(x_{1}),\mu(x_{2}),...,\mu(x_{n})]^{T},\sigma^{2}\bm{R})$ (5)
Where $\mathcal{MN}(\cdot)$ is a multivariate normal distribution,$y_{x,y}(\cdot)$ is the real-valued QoI at a point location with coordinates $(x,y)$, $\bm{x^{d}}\in\mathcal{X}$ is the input vector to the simulation, $\mu(\cdot)$ is the mean function, $\sigma^{2}$ is the unknown variance, and $\bm{R}$ is the correlation matrix. The mean function $\mu(\cdot)$ of a GP is typically modeled as a regression:
$\mu(\bm{x})=\sum_{i=1}^{q}h_{i}(\bm{x})\theta_{i}$ (6)
Where $h_{i}(\cdot)$ are the mean basis functions and $\theta_{i}$ the corresponding regression coefficient. For computational efficiency, the basis functions $h_{i}(\cdot)$ are common between points that belong to the same QoI and only the coefficients $\theta_{i}$ can vary. The elements of the correlation matrix $\bm{R}$ in (Eq. 5), are the values of a chosen correlation function between observation vectors $\bm{x_{i}},\bm{x_{j}}$:
$c(\bm{x_{i}},\bm{x_{j}})=\Pi_{k=1}^{d}c_{k}(x_{ik},x_{jk})$ (7)
Where $c_{k}$ is the output of the correlation function for the $k^{th}$ coordinate of the two input vectors. There are many options for correlation functions and here we use the rational quadratic kernel, which is created by adding squared exponential kernels across different length scales because it would capture some of the multiscale effects of the DNS. Other kernels failed, as discussed later. Overall, the unknown parameters in the GP formulation include the regression coefficients $\theta$, the variance $\sigma^{2}$, and the range parameters from the kernel as a vector $\bm{\gamma}$. In this formulation, the marginal likelihood function after integrating out ($\theta,\sigma^{2}$) becomes:
$\mathcal{L}(y^{d}|\bm{\gamma})\propto|\bm{R}|^{-1/2}|\bm{h}^{T}(\bm{x}^{d})\bm{R}^{-1}\bm{h}^{T}(\bm{x}^{d})|^{-1/2}(S^{2})^{-(\frac{n-q}{2})}$ (8)
Where $S^{2}=(\bm{y^{d}})^{T}\bm{Q}\bm{y^{d}}$, $\bm{Q}=\bm{R}^{-1}\bm{P}$, and $\bm{P}=\bm{I_{n}}-\bm{h}(\bm{x^{d}})[\bm{h}^{T}(\bm{x^{d}})\bm{R}^{-1}\bm{h}(\bm{x^{d}})]^{-1}\bm{h}^{T}(\bm{x^{d}})\bm{R}^{-1}$. $\bm{I_{n}}$ is the identity matrix of size $n=64$, which is equal to the number of simulation runs. The range parameters $\gamma$ are then estimated by the modes of the marginal joint posterior distribution:
$\hat{\bm{\gamma}}=argmax_{(\gamma_{1},...,\gamma_{p})}\mathcal{L}(y^{d}|\bm{\gamma})\pi(\gamma_{1},...,\gamma_{p})$ (9)
To train the GP, we use n design points from the DNS ($\bm{x^{d}}=[\bm{x_{1}^{d}},\bm{x_{2}^{d}},...,\bm{x_{n}^{d}}]$) and the output is the QoI at these points $\tilde{\dot{r}}(\bm{x^{d}})$ for the parameter estimation described above. After training, the GP can be used for predicting the QoI at a specific position (x,y) on the slab burner boundary from a set of new input $\bm{x^{*}}$.
3.2 Hierarchical Multiscale Surrogate (HMS)
Figure 2: Behavior of the basis functions at increasing scales. Figure from [51].
Hierarchical and multiscale models typically have an identified hierarchy of approximations and make joint inference on the data by combining these model components in some rational fashion [52, 53]. Hierarchical models also have parallels to deep learning models which implement sequential function compositions to learn a data generation mechanism [54], multiresolution Analysis (MRA) implementing wavelet decompositions, ([55]), geometric analysis ([56, 57]) involving diffusion maps to analyze datasets at different scales and Hierarchical Radial Basis Functions (HRBF) ([58, 59, 60]). The combination of basis functions at different scales to produce a sparse representation is the core idea in most of these [52] though as we show in recent work [51], it is often efficient in the prediction phase to use only an optimal scale rather than a mix of scales. Hierarchical approximations also find applications in data visualization [61]. Recent work on explaining the success of deep learning has attributed it to the ability of these networks to accomplish hierarchical learning efficiently [62]. This approach for sparse representation, introduces a scale parameter $s$ and defines a mapping between $s$ and the corresponding approximation space $\mathcal{H}_{s}$ in a hierarchy of spaces. The idea of exploiting the inherent correlation structure in the data at multiple levels follows directly from [63, 64]. The hierarchical models are able to analyze irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting using scale and data dependent basis functions $B^{s}$, constructed by sampling suitable kernels at successively finer scales. Basis selection to approximate the observed data efficiently [40] uses a greedy scheme with forward selection and backward deletion phases that provides good performance avoiding the performance bottleneck of other strategies [51].
Figure 3: Hyperparameter tuning for the HMS surrogate. $\delta$ describes the upper bound tolerance for the current scale tolerance ($\epsilon_{s}<\delta$), $s$ is the number of scales, and $k=[1,5]$ refers to a training data fold. A vertical identifier is placed at the scale where the MSE is $<2e^{-05}$. Smaller compression ratios are better for fast prediction. Based on these results, we select $s=15,\delta=0.01$ for the two hyperparameters.
Input: $D=(X,y)$, $\omega$, $\epsilon_{0}$ Initialize: $s=0$, $B^{s}=[\cdot]$, $\Theta^{s}=[\cdot]$, $C^{s}=[\cdot]$, $Af^{s-1}=0$, $t_{s}=y$
While $s\leq\omega$ do:
Step 1. $B_{s},\Theta_{s},C_{s},\mathcal{E}\leftarrow Forward\_selection(\epsilon_{s},D,s,t_{s})$
Step 2. $B_{s},\Theta_{s},C_{s}\leftarrow Backward\_deletion(\epsilon_{s},B_{s},\Theta_{s},C_{s},\mathcal{E},t_{s})$
Step 3. $B^{s},\Theta^{s},C^{s}\leftarrow Append(B_{s},\Theta_{s},C_{s})$
Step 4. Current scale approximation: $f_{s}\leftarrow B_{s}\Theta_{s}$
Step 5. Current overall approximation: $Af^{s}\leftarrow Af^{s-1}+f_{s}$
Step 6. Update scale: $s\leftarrow s+1$
Step 7. Update target for next scale: $t_{s}\leftarrow t_{s-1}-f_{s-1}$
Step 8. Update tolerance: $\epsilon_{s}$
End While
where
$D$: Dataset (X,y), such that $X\in\mathbb{R}^{n\times d}$ and $y\in\mathbb{R}^{n}$, $s$: scale number, $\omega$: truncation scale, $\epsilon_{0}$: tolerance at scale 0, $Af^{s}$: final approximation till scale $s$, $t_{s}$: target at scale $s$, $B_{s},\Theta_{s},C_{s}$: basis set, corresponding weights, and sparse representation at scale $s$, $B^{s},\Theta^{s},C^{s}$: cumulative basis set, corresponding weights, and sparse representation till scale $s$
Algorithm 1 The Hierarchical Multiscale Approach [51]
Algorithm 1 summarizes the key steps. A multiscale kernel (Eq. 10) [51], which uniquely defines a customized associated approximation space (RKHS) is used. s: scale of function, T: size of domain.
$K(x,y)=\sum_{s\in\mathcal{I}}\zeta_{s}\sum_{j\in r_{s}}\psi_{j}^{s}(x)\psi_{j}^{s}(y),\quad\psi^{s}_{j}(z)=\exp\Bigg{(}-\frac{||z-z_{j}||^{2}}{\kappa_{s}}\Bigg{)},\quad\kappa_{s}=\frac{T}{2^{s}}$ (10)
At each scale for a given target $t_{s}$, the following optimization problem is solved
$\min_{\beta_{s}}||t_{s}-K_{s}\beta_{s}||_{2},\quad\text{subject to }||\beta_{s}||_{0}\leq p$
Here $K_{s}=[b_{s}^{j}]$ is the full set of n functions at scale $s,p$ are some positive constant and $||\beta_{s}||_{0}$ counts the number of non-zero values in $\beta_{s}$. Since, this is a non-convex problem, an approximation of the problem corresponding to $||\cdot||_{1}$ norm is obtained. The greedy forward selection starts with an empty basis set $B_{s}$, and recursively adds functions from the set $K_{s}$ which provide the highest reduction in the current residual. This is continued until the tolerance $\epsilon_{s}$ is satisfied by the residual $r^{t}$ as $\frac{|r^{T}b^{j}_{s}|}{||b_{s}^{j}||_{2}^{2}}<\epsilon_{s}$. Before training the HMS surrogate, we performed hyperparameter tuning to select the number of scales and tolerance bounds. The tuning process is about the trade-off between approximation error and the compression ratio of the data $|C^{s}|/|X|$. As more scales $s$ are added, the training error reduces, while the need for more data in the approximation increases. The selected number of scales and tolerance are selected such that we achieve an acceptable order of error magnitude ($\sim e^{-05})$, while also maintaining the compression ratio as low as possible for fast prediction. Based on our test runs (see Fig. 3 for results from best runs), we selected $s=15$ scales and the tolerance bound as $\epsilon_{s}<\delta=0.01$.
3.3 Uncertain Inputs and Ensemble Strategy
The relevant input quantities to ABLATE are summarized in Table 1. To create accurate surrogates, the simulation ensemble needs to be selected such that it covers the domain and codomain of the simulation as best as possible. Some of the input quantities are considered uncertain and therefore have to be sampled for the ensemble. For the ensemble runs, the geometry remains fixed because the simulation runs for a relatively small real time (on average 25-75ms). We do not expect the geometry change to be significant during this short time given that the maximum regression rate of PMMA experimentally is measured up to 0.15mm/s [65], which justifies our assumption for the fixed geometry during a short simulation. We initially considered all possible reaction rate parameters of the Arrhenius expression (Eq. 11) and used local and global sensitivity analysis to select the reaction rate parameters to be included in the simulation ensembles, as described in Section 3.3.1. Only the selected reaction parameters are sampled for the ensembles, the remaining ones are fixed at expected values from literature. The inlet velocity factors are selected such that the ABLATE simulation corresponds to reasonable oxidizer flux $G$ of an equivalent slab burner experiment [5], so we selected velocities that correspond to $G{\in}[5,20]\frac{kg}{m^{2}s}$. Of the fuel parameters $\theta_{f}$, the latent heat of sublimation $l_{v}$ is considered uncertain and sampled given an estimated range from expert input, but is later calibrated with available experimental data. To generate the 64 simulation ensemble, the uncertain input quantities were selected with Latin Hypercube Sampling (LHS), where each sample is the only one in each axis-aligned hyperplane containing it. The values for the uncertain inputs of each simulation of the ensemble are shown in Table LABEL:tab:ensemble1 and LABEL:tab:ensemble2 in the Appendix.
Table 1: Relevant input quantities for the ABLATE coupled flow/combustion solver.
Quantity (units) Symbol Type Description/Notes
Velocity factor (m/s) $V_{in}$ Uncertain The inlet velocity profile is assumed parabolic
Activation Energies (cal/mol) $E_{*}$ Uncertain The energy required to form the transition state for a given reaction
Temperature exponents (-) $b_{*}$ Uncertain See Eq. 11
Pre-exponential factors (cm3 mol-1 cal -1) $A_{*}$ Uncertain Empirical constant between temperature and rate coefficient
Geometry $\mathcal{L}_{slab}$ Fixed -
Latent heat of sublimation (J/kg) $l_{v}$ Uncertain The amount of heat energy required for PMMA to sublimate (fuel param.)
Sublimation temperature (K) $T_{f}$ Fixed Temperature at which PMMA sublimates (fuel param.)
3.3.1 Feature selection of Reaction Rate Parameters
High fidelity models of fuel combustion typically involve a significant number of species and reactions. Despite the large dimensionality of the reaction rate parameter space, it has been observed that in many combustion systems, just a few parameters often have a more significant impact on the uncertainty of chemical model outputs compared to others [66, 67, 68]. Based on the literature, we use sensitivity analyses and feature selection to systematically reduce the parameter space. The result is a reduced set of reaction rate parameters that we consider uncertain and include as features in the ensembles that train the surrogate models. The local sensitivity analysis (LSA) investigates the impact of small variations around a nominal parameter value on some reaction output (ignition delay). LSA employs a truncated Taylor expansion to approximate the sensitivity coefficient to a first-order level, which is advantageous due to its simplicity and fast computation speed, making it a valuable tool for gaining qualitative insights into the model behavior. Global uncertainty analysis (GSA) is a statistical approach and enables us to determine how uncertainties in ignition delay can be attributed to uncertainties in model parameters. The Sobol method [69] is a variance-based global sensitivity analysis (VGSA) that ranks the importance of parameters by quantifying the extent to which the conditional variance caused by a parameter explains the variance in the model output. The Sobol method evaluates total sensitivity indices for each input parameter, providing estimates of the parameter’s effect and its interactions with other parameters on the variation of ignition delay. Feature selection is a process to select a smaller set of features than those available and can be accomplished in many ways depending on the goal [70]. By leveraging these methods to reduce the uncertain parameter space in chemical models we do the following, in order:
  1. 1.
    Local sensitivity analysis to identify most influential reactions out of all reactions in the mechanism.
  2. 2.
    Global sensitivity analysis on the parameters of the reactions identified in the previous step to find the most important rate constant parameters contributing to the chemical model output.
  3. 3.
    Feature selection between the reaction rate parameters with the objective being to minimize the testing MSE of the HMS surrogate. The other slab burner covariates ($x,y,l_{v},V_{in}$) all remain in the surrogate.
We have implemented this framework to identify sensitive reactions and parameters of a reduced MMA mechanism, consisting of 67 species and 263 reactions. The dependent variable the sensitivity analyses are completed with is the ignition delay time (IDT), due to its significance in combustion kinetics [71]. The first-order sensitivity index is the change of MMA mass fraction with respect to the rate constants ($k_{i}$) of the $i$th reaction, e.g., $\dfrac{\partial\ln(MMA)}{\partial\ln(k_{i})}$, and is evaluated at the pressure of 1 atm and over a range of initial temperature of reactor between 600K and 2000K. Typically, forward rate constants ($k_{i}$) of the MMA mechanism are calculated via the 3-parameter Arrhenius expression:
$k_{i}(T)=A_{i}T^{\beta_{i}}\exp\left(-\frac{E_{i}}{RT}\right),$ (11)
where $T$ is the temperature, and $E_{i}$, $A_{i}$, and $\beta_{i}$ are the activation energy, pre-exponential factor, and temperature exponent coefficient of the $i$th reaction, respectively. By identifying the sensitive reactions via local sensitivity analysis, it becomes computationally feasible to perform VSA on their associated associated rate parameters ($A_{i},b_{i},E_{i}$). To conduct VGSA, we sample all parameters from a uniform probability distribution function (PDF) and compute the total Sobol sensitivity index ${S_{T}}_{i}=\frac{\mathbb{E}[Var(IDT_{i}|k_{\sim i})]}{Var(IDT)}$. However, in the combustion literature, it is known that there are multiple sources of correlation among these parameters [72] [73]. For our VSA, we consider a 20% uncertainty bound around the nominal values of the Arrhenius parameters associated with the most influential reactions and use TChem [74, 75, 76, 77, 78, 79] solvers for the sensitivity analyses. TChem computes the ignition delay time for a 0-dimensional constant volume homogeneous gas reactor. TChem defines IDT as the time when the build-up of $\mathrm{OH}$ radicals peaks. From the LSA, the most important reactions are the following:
$\mathrm{Reaction\hskip 2.84526pt\#1:\hskip 2.84526ptH+O_{2}\Leftrightarrow O+OH}$
$\mathrm{Reaction\hskip 2.84526pt\#257:\hskip 2.84526ptMMA\Leftrightarrow T-C_{3}H_{5}+CH_{3}OCO}$
$\mathrm{Reaction\hskip 2.84526pt\#250:\hskip 2.84526ptMMA+CH_{3}\Leftrightarrow MJ+CH_{4}}$
$\mathrm{Reaction\hskip 2.84526pt\#249:\hskip 2.84526ptMMA+H\Leftrightarrow PJ+H_{2}}$
Figure 4 illustrates the results from the VGSA. The total sensitivity index is evaluated using 6000 samples drawn from the parameter distributions using LHS, using the Monte -Carlo (MC) estimator proposed by [80] and implemented in the parallel object-oriented library DAKOTA [81]. The figure shows that the activation energy in reaction 257 exhibits the most impact on ignition delay. This particular MMA reaction corresponds to chain branching highly reactive species, which play a crucial role in sustaining the combustion process by breaking down and producing radicals. Other important parameters include the activation energies of reaction 1 and 250 as well as the temperature exponent for reaction 249. Furthermore, we observed the rest of the parameters ($A250,b257,A1,E249,A249,A257$) to have a negligible effect on the IDT. Consequently, these parameters can be considered deterministic when producing simulation samples to train a surrogate model, as their uncertainties do not significantly impact the overall variability in the IDT.
Figure 4: Left: Effect of the most important Arrhenius parameters on ignition delay for our MMA mechanism. Right: Total sensitivity indices of the Arrhenius parameters of the four influential reactions identified by local sensitivity analysis. The 4 most important parameters ($E257,b249,E1,E250$) are considered most important and are therefore part of the uncertain simulation ensemble. The rest of the parameters are considered deterministic and are fixed, since they have minimal impact on ignition delay.
Figure 5: Surrogate feature selection between reaction rate parameters. The dashed red line indicates the minimum MSE. $E1$ and $b249$ achieve similar results. The test MSE is added over all testing folds. We note that adding all 4 of the reaction rate parameters results in an increase to testing MSE by an order of magnitude. Based on these results, we select $b249$ as the only reaction rate to be included in the surrogate.
To complete the feature selection, the testing MSE of the HMS surrogate is evaluated across different testing folds covering the entire dataset. Each testing fold is predicted with different models arising from the different combinations of reaction rate parameters and all the other slab burner predictor features ($x,y,l_{v},V_{in}$). The testing MSE evaluated here is an addition of the error at each fold. The results of the feature selection are shown in Fig. 5. Based on our analysis, we only include $b249$ as the only reaction rate feature in the surrogate models. While it may be unexpected that $b249$ alone produces the best surrogate model for generalization even if it is not the most sensitive parameter to ignition delay, we note that these parameters are not entirely independent of each other. Therefore, the sensitivity analysis (which is conducted with ignition delay as the dependent variable) is only a part of the selection process, whereas directly testing the surrogate model against the slab burner QoI (the regression rate) is ultimately the deciding factor, as regression rate is the quantity we want to predict.
3.4 Forward Uncertainty Propagation
After the development of the surrogates, they can be used to propagate forward the uncertainty in the inputs $\bm{x}=[x,y,b_{249},l_{v},V_{in}]$ to the QoI $\tilde{\dot{r}}$. The goal is to approximate the probability distribution of the regression rate and its statistics, which can be achieved via MC sampling methods [82] (Alg. 2). Forward UQ is possible given that the surrogates are fast to sample. To know when the MC sampling has converged, we monitor the mean and variance of the posterior distribution for the regression rate and stop when those its statistics difference are within a tolerance of $\epsilon_{*}=e^{-5}$. For the inputs, the velocity factor is assumed uniform throughout the parameter range. The fuel and chemistry parameters $l_{v},b_{249}$, are assumed normal with the mean being the nominal default DNS value and a variance that results in the distribution covering values ±$\pm 25\%$ from that nominal value. Given that the regression rate is highly dependent on the location of interest as known from our previous work with experiments, we consider the results in different regions of the slab burner. Experimentally, the front part ($x\leq 40$mm) is expected to achieve the highest regression due to the ramp effects and ignition location, the middle part ($40<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects. In each case, we uniformly sample the possible boundary locations.
Given a portion of the slab (front, middle, or rear) with locations $(x,y)\in\mathcal{S}_{B}$
Assume prior distributions for:
$V_{in}\sim\mathcal{U}(V_{min}=3.917\frac{m}{s},V_{max}=15.398\frac{m}{s})$
$b_{249}\sim\mathcal{N}(2.54,0.029)(-)$
$l_{v}\sim\mathcal{N}(840890,4.91\times 10^{9})\frac{J}{kg}$
[…]
While $\|\sigma_{\bm{y^{*}}}^{2,new}-\sigma_{\bm{y^{*}}}^{2,old}\|\geq\epsilon_{\sigma^{2}}$ AND $\|\mu_{\bm{y^{*}}}^{new}-\mu_{\bm{y^{*}}}^{old}\|\geq\epsilon_{\mu}$ do:
Step 1. Draw a sample from each of the input quantities to create the new design vector $\bm{x^{*}}=[x,y,E^{*}_{1},l_{v}^{*},V_{in}^{*}]$
Step 2. Get the new prediction for the regression rate from the surrogate: $\bm{y^{*}}=\tilde{\dot{r}}(\bm{x^{*}})$
Step 3. Append corresponding probability distribution $p({\tilde{\dot{r}}})=\mathrel{+}\bm{y^{*}}$
Step 4. Update statistics $\sigma_{\bm{y^{*}}}^{2,new},\sigma_{\bm{y^{*}}}^{2,old},\mu_{\bm{y^{*}}}^{new},\mu_{\bm{y^{*}}}^{old}$
End While
Algorithm 2 Forward UQ: Estimation of the distribution of the regression rate using the surrogate
3.5 Model Calibration under Uncertainty
Finally, we demonstrate the Bayesian calibration of two key parameters: the latent heat of sublimation ($l_{v}$)–a fuel parameter in the DNS model—and the temperature exponent for reaction 249 ($b_{249}$)–a parameter in the chemical reaction mechanism. Both parameters are calibrated using regression rate measurements obtained from slab burner experiments. The calibration of physics-based models necessitates addressing uncertainties arising from incomplete and noisy experimental data, as well as the model’s limitations in capturing complex physical phenomena. Traditional deterministic inverse methods are insufficient for quantifying these uncertainties. Therefore, we adopt Bayesian approaches for statistical inverse analysis, as outlined in [83, 84, 85], which provide a probabilistic framework for UQ in model calibration. The solution to the Bayesian calibration problem is the probability distribution functions (PDFs) of the parameters ($\bm{x}_{calib}={l_{v},b_{249}}$) given the experimental data $\mathbf{D}$, such that
$\pi_{post}(\bm{x}_{calib}|\mathbf{D})=\frac{\pi_{like}(\mathbf{D}|\bm{x}_{calib})\pi_{prior}(\bm{x}_{calib})}{\pi_{evid}(\mathbf{D})},$ (12)
where the prior $\pi_{prior}(\bm{x}_{calib})$ represents initial knowledge about parameters before observing the data, likelihood $\pi_{like}(\mathbf{D}|\bm{x}_{calib})$ is derived from a noise model representing data and model uncertainty, and evidence PDF $\pi_{evid}(\mathbf{D})$ is a normalization factor ensuring the posterior $\pi_{post}(\bm{x}_{calib}|\mathbf{D})$ is a PDF.
The primary computational challenges in the Bayesian calibration of coupled flow and combustion DNS models like ABLATE include the high computational cost of each forward run and the extensive parameter space, which makes sampling-based Bayesian algorithms like Markov Chain Monte Carlo (MCMC) computationally prohibitive. To address these challenges, we use the HMS surrogate model that substantially reduces the computational cost of mapping inputs to regression rate, i.e., $\tilde{\dot{r}}=\mathbf{Y}(\bm{x}_{calib})$ within the Bayesian solution. Assuming an additive noise model and Gaussian random variables for the data noise with zero mean, we postulate the following form for the likelihood:
$\ln\left(\pi_{like}(\mathbf{D}|\bm{x}_{calib})\right)\propto\left[\mathbf{Y}(\bm{x}_{calib})-\mathbf{D}\right]^{T}\bm{\Gamma}^{-1}\left[\mathbf{Y}(\bm{x}_{calib})-\mathbf{D}\right],$ (13)
where $\bm{\Gamma}$ is a diagonal covariance matrix encapsulating the noise at each data point. The process to measure the regression rate from image data is outlined in [5] and includes capturing snapshots of the PMMA solid fuel profile as it regresses during the burn using high-speed cameras. The resulting profiles are then either manually traced and segmented or are segmented using machine learning [6]. The data used for this work were all manually traced. By knowing the height of the fuel profile over time and at each location, the regression rate on the boundary can be estimated by averaging the height difference over the observed experimental time of the snapshots. The available experiment data are shown in Fig. 6, taken for an experiment with oxidizer flux $G=8.08\frac{kg}{m^{2}s}$.
Figure 6: Available regression rate measurements from an experiment with oxidizer flux $G=8.08\frac{kg}{m^{2}s}(\equiv V_{in}=6.258m/s)$.
We note that we have taken many steps to guarantee the closest possible equivalence between the regression rate from the surrogate and the experimental data, although there are a few differences that need to be disclosed. The simulation data (which the surrogate is trained with) are estimates of the regression from the heat flux balance at the slab burner boundary and are computed over short simulation times (typically a few tens of milliseconds). To the contrary, experimental regression rate is purely based on the fuel profiles, carries uncertainty due to the manual tracing, and the snapshots used to compute the regression rate estimates span over a few seconds due to waiting for the fuel specimen temperature to get to steady-state. With such data available, the steps we have taken to guarantee the closest possible equivalence between the two sets of data are the following: a) remove time-dependence on the regression rate by time-averaging either the profile difference (experiments) or the direct regression rate estimate (simulation), b) have the simulation use an isothermal boundary condition to best approximate the steady-state conditions of the experiment data collection, and c) estimate the variance in the experimental data used in the Bayesian calibration from 5 repetitions of the manual tracing of the slab profiles (see Fig. 6).
4 Results and Discussion
This section provides the results after implementation of the methods introduced in Section 3. The section starts with a comparison via cross validation between the two surrogates GP and HMS. Then, we use HMS for forward UQ to estimate the probability distribution of the regression rate from the uncertain inputs in 3 regions of the slab burner: front, middle, and rear. Lastly, we demonstrate Bayesian inference of the fuel parameter latent heat of sublimation ($l_{v}$) and the temperature exponent of reaction 259 ($b_{249}$) with experimental slab burner regression rate data.
4.1 GP vs. HMS comparison
To compare the two surrogate models, we first evaluate their training and testing accuracy on the available dataset via 5-fold cross validation [86]. We split the dataset of the 64 DNS simulations in 5 folds, 4 of which include the 13/64 simulations as part of the testing set and the remaining 51/64 simulations as training. The last fold includes 12/64 simulations as testing and the remaining 52/64 simulations as training. When a simulation is picked to be part of a training/testing set, all the boundary locations are included with it, we did not sample based on location. With the cross-validation approach, the goal is to evaluate the resulting surrogate models on the entire available dataset and particularly their ability to generalize well and predict correctly out-of-sample, i.e., from data they have not been trained on.
Fig. 7 shows some of the results from the first fold for both GP and HMS. Both surrogates are trained correctly and predict the training set perfectly. One obvious difference between the two models is that GPs inherently predict with a standard deviation which can be translated to a $95\%$ confidence interval, unlike HMS. Therefore, a GP may be more practical as evident in some of the results shown in Fig. 7 (e.g. see $IDs=8,12,13$), since it is able to capture all the simulation output within the confidence interval even if the mean is not precise. Comparing the GP mean with HMS, we conclude that both models appear to have similar performance out of sample, but the HMS follows the simulation data visibly closer when non-linear effects are present (e.g., see $IDs=2,11$). We also conclude that neither model is able to fully capture non-linearities as evident in e.g., $IDs=10,9,11$. Further investigation into this issue brought forward the following conclusions:
  1. 1.
    The GP performance is sensitive to the selection of kernel. Rational quadratic is the only kernel that produced a correct GP trained on the DNS data. The rational quadratic is a collection of exponential kernels, which captures some of the multiscale effects of the problem and justifies why the GP performs similarly to HMS. See Fig. 8(a) for an example of a poorly fit GP with Matérn kernel. The HMS on the other hand adaptively selects a set of basis functions that work for a large class of problems without heuristic choices of kernel.
  2. 2.
    Even though the DNS is solving a multiscale problem, both surrogate models are unable to capture all of the scales since the inputs are all far field inputs (far field velocity, parameters). For example, in Fig. 7, the non-linearities at the end of $ID=11$ are missed, and HMS is closer to the top of the regression rate spike than the GP. The models are attempting prediction of a boundary quantity that is dependent on the combustion process, but none of the far field inputs they are trained on fully carry this multiscale information. There is no inherent structure in the surrogates to be able to estimate the multiscale combustion process, unlike the DNS that fully solves the system.
  3. 3.
    As a demonstration, when a combustion-related quantity ($\dot{q}_{Net}$) is added as an input to the surrogate, the HMS fully captures the multiscale effects and is able to predict out of sample with precision as shown in Fig. 8(b). We note however that this would not be a useful surrogate in practice, as it is impossible to know the heat flux a priori without solving the combustion problem, which defeats the purpose of having a surrogate.
We show some more comparisons between the two models from the other cross-validation folds in the Appendix in Fig. LABEL:fig:HMSvGP_2.
Fig. 9 shows the training and testing performance of the two models across all the data folds, measured as a normalized percentage error from the true value of the prediction. The GP, on average, is expected to achieve an error of $1.59\times 10^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Overall, our conclusion is that both of these models are acceptable for usage in UQ or otherwise. The fact that both of these models are able to achieve an error $<20\%$ when predicting multiscale boundary quantities just from a few far field inputs is significant, particulary considering that they take a few $ms$ to run compared to 24 hours for the DNS.
Figure 7: Comparison of GP surrogate and HMS for select simulation IDs belonging to the first fold of the cross validation. Figure titles mark whether the data were in the train or test set. As expected, both surrogates predict the training set perfectly. The GP inherently includes a standard deviation in the predictions which is translated to a $95\%$ confidence interval, adequately including the simulation data within these bounds. Both models seem to have similar mean predictions. When the regression rate includes nonlinear effects (e.g., spike at $70mm$ for $ID=2$ or spike at $80mm$ for $ID=11$ ), the HMS mean follows the trend visibly a bit closer than the GP.
(a) Fig. 8a     Example of a GP trained on the DNS data with Matérn kernel. The optimizer has not converged as evident by the very large variance bounds and the poor fit of the mean. Rational quadratic is the only kernel that produced a correct GP trained on the DNS data.
(b) Fig. 8b     Example of a HMS trained on the DNS data including the net radiation at the boundary ($\dot{q}_{Net}$). The HMS fully captures the multiscale effects and is able to predict out of sample with precision.
Figure 9: Performance of GP and HMS across all data folds of the cross validation. The GP, on average, is expected to achieve an error of $1.59\times 10^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Overall, our conclusion is that both of these models are acceptable for usage in UQ or otherwise.
4.2 Forward UQ with HMS
In this section, we demonstrate the use of the HMS for propagating uncertainty from the inputs to the regression rate as shown in Alg. 2. Approximating the distribution of the regression rate from the uncertain input space is useful for decision making and in this case to be able to bound the expected performance of the slab burner in terms of regression rate. We consider 3 areas of study for the slab burner since the regression rate is highly dependent on the location. Experimentally, the front part ($x\leq 40$mm) of the slab is expected to achieve the highest regression due to the ramp effects, the middle part ($40<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects from the oxidizer flow. We do not expect the surrogate models to be able to match the experimental result however, as the 2D DNS does not either as shown previously in Fig. 7 where regression is low in the front part. In each forward propagation, we select boundary locations uniformly that adhere to the restrictions of the location. The velocity factor $V_{in}$ is also uniformly sampled and its range corresponds to an oxidizer flux of $G_{min}=5\frac{kg}{m^{s}}$ and $G_{max}=20\frac{kg}{m^{2}s}$, which are appropriate for this slab burner. The remaining two parameters are the latent heat of sublimation $l_{v}$ and the temperature exponent for reaction 249 ($b_{249}$) which are assumed normal with the mean being the nominal default value of the DNS and a variance that corresponds to a distribution that covers ±$\pm 25\%$ above or below the nominal value. The forward propagation uses traditional Monte-Carlo sampling and the results for the distributions of the regression rate as shown in Fig. 13. In the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.022$ mm/s, in the middle of the slab $\mu_{m}=0.081$ mm/s, and in the rear $\mu_{b}=0.147$ mm/s. The front and rear distributions appear to have a more central tendency compared to the middle part where the density is more evenly distributed on a wider range of values.
Figure 10: $\mu_{f}=0.022,\sigma_{f}=0.021$mm/s
Figure 11: $\mu_{m}=0.081,\sigma_{m}=0.039$mm/s
Figure 12: $\mu_{b}=0.147,\sigma_{b}=0.029$mm/s
Figure 13: Forward propagation of uncertainty to the regression rate from uncertain inputs at 3 locations of the slab burner corresponding to $G=[5,20]\frac{kg}{m^{2}s}$. From left to right: $x_{slab}\leq 40$mm , $40<x_{slab}\leq 70$mm, $x_{slab}>70$mm. The 2D DNS estimates the regression to be lowest in the front and largest in the back of the slab. From experiments we know the front achieves the highest regression rate which is not captured by the DNS.
4.3 Bayesian model calibration using regression rate measurement data
In this section, we use the HMS surrogate model in conjunction with available experimental data (Fig. 6) to calibrate two key physical parameters of the DNS model: the latent heat of sublimation $l_{v}$ and the temperature exponent $b_{249}$. As described in Section 3.5, we apply Bayesian inference to calibrate these parameters while quantifying the associated uncertainty in model parameters. The prior distributions for the parameters are the same as used for the forward UQ, namely $b_{249}\sim\mathcal{N}(2.54,0.029)(-)$ and $l_{v}\sim\mathcal{N}(840890,4.91\times 10^{9})\frac{J}{kg}$. They are assumed normally distributed with the mean being the nominal default DNS value and a variance that results in the distribution covering values ±$\pm 25\%$ from that nominal value. The mean is taken from expert input and existing literature. To solve the Bayesian calibration problem, we employ the Random Walk Metropolis algorithm [87], a Markov Chain Monte Carlo (MCMC) method used to draw samples from the posterior distribution. To thoroughly explore the posterior distributions, we run 5 independent chains of 10,000 iterations each, starting from different initial points within the prior: both parameters at the edges of their feasible domain, both at their mean values, or one at the minimum and one at the maximum of their respective domains. The first 10% of each chain is discarded as “burn-in,” which is standard practice to allow the chain to converge towards the true posterior before sampling begins. Figs. 14, 15 present trace plots for one representative MCMC chain initialized with both parameters at the minimum of the prior range. Additionally, the autocorrelation plot (Fig. 17) is provided as a diagnostic tool to assess the independence of the samples. As shown, samples spaced 60 iterations apart exhibit minimal correlation, indicating that the MCMC chain is progressing efficiently and producing low-correlated, independent samples, and that the chain has appropriately explored the posterior distribution.
To arrive at the final marginal posterior estimates (see Figs. 18, 19), we combine all 50000 samples from all 5 chains. A main conclusion from the results is that the default value for $l_{v}$ at the default value of 840890 J/kg is too large to allow the surrogate to achieve the experimental results in the front of the slab. Instead, the value should be lower ($\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$). The resulting posterior distribution for $b_{249}$ ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) has a mean closer to the default value 2.54 from the DNS.
Figure 14: Bayesian calibration results: chain for $l_{v}$ starting with parameter at its prior minimum value.
Figure 15: Bayesian calibration results: chain for $b_{249}$ starting with parameter at its prior minimum value.
Figure 16: Bayesian calibration results: scatter plot of the chain for $l_{v}$ and $b_{249}$.
Figure 17: Bayesian calibration results: autocorrelation plot of the chain.
Figure 18: Final marginal posterior distribution of $l_{v}$ using HMS and experimental data. We note that the posterior is far from the nominal value of 840890 $J/kg$ that the DNS uses as default. The calibration shows that $l_{v}$ should be lower ($\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$) for the HMS to get results closer to the data in the front part of the slab.
Figure 19: Final marginal posterior distribution of $b_{249}$ using HMS and experimental data. The statistics of the distribution at ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) are close to the default value of 2.54 used by the DNS.
To further demonstrate the impact of calibration, we show in Fig. 20, a forward UQ in two cases: “calibrated” and “not calibrated” for a specific oxidizer flux of $G=8.08\frac{kg}{m^{2}s}$, which is the same conditions used in calibration. The “not calibrated” case includes a forward UQ process with the same priors listed in Alg. 2. Instead, the “calibrated” case takes the prior distribution for $l_{v}$ and $b_{249}$ to be their posteriors from the Bayesian calibration. The result is a shift of the probability density for the regression rate towards higher values in all parts of the slab. The result verifies our previous observation: the DNS and subsequently the surrogates underestimate the regression rate in the front part of the slab, when used with the default values for the parameters in forward UQ. However, when we use calibrated priors for the parameters, we find the surrogate to produce samples for the front (now up to 0.14mm/s) that match the experiment. That also results in overestimating the regression rate in the middle and rear parts of the slab. The two calibrated parameters basically have to compensate for the model inadequacy between the 2D DNS and the experimental data and their effect is not sufficient to make the surrogate be simultaneously accurate in all slab locations. When uncalibrated, the surrogate underestimates the regression rate in the front part, correctly predicts the middle, and overestimates the rear part of the slab. When calibrated priors are used for the parameters, the surrogate correctly predicts the front, but further overestimates the middle and rear part of the slab. Overall, the conclusion is that there is not one set of ($l_{v},b_{249}$) that result in correctly predicting the regression rate in all locations of the slab. Calibration helps to identify the correct ranges to be used for specific locations. A next step would then be to have an accurate surrogate for different locations or conditions by simulating the DNS within the corresponding bounded parameter regimes.
Front slab
Middle slab
Rear slab
Figure 20: Forward UQ for a fixed velocity corresponding to $G=8.08\frac{kg}{m^{2}s}$ with default parameter priors and calibrated priors. The surrogate underestimates the regression rate in the front part of the slab, when used with the default values for the parameters in forward UQ. When we use calibrated priors for the parameters, we find the surrogate to produce samples for the front that match the experiment better.
5 Conclusion and Future Work
The main goal of this paper was to systematically address challenges related to all aspects of performing a complete UQ analysis of a complicated physics-based simulation like a 2D slab burner DNS. We provided insights related to the development of an acceptable surrogate model, the propagation of uncertainty of the DNS inputs to the QoI, and the calibration of DNS parameters based on observations of experimental data as well as interpreting these results for further DNS development. The DNS we used in this work was ABLATE (Ablative Boundary Layers At The Exascale), an in-house DNS solver developed for simulating a multiphase 2D slab burner setup.
First, we generated an ensemble of 64 simulations that was informed by a series of sensitivity analyses to identify the most important reaction rate parameters. The simulation data was then used to develop two surrogate models: a Gaussian Process (GP) and a hyperparameter tuned Hierarchical Multiscale Surrogate (HMS). Our expectation was that the HMS would outperform the GP because of the multiscale effects of combustion. We tested both models under cross-validation for their ability to predict out of sample in parameter spaces they have not been trained on. We found that, overall, they performed very closely. The GP, on average, is expected to achieve an error of $1.59e^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Our conclusion is that both of these surrogate models are acceptable for usage in UQ or otherwise. We do note however that the GP performance is sensitive to the choice of kernel and specifically kernels that cannot handle multiple scales (such as the rational quadratic) will result in poor models. HMS demonstrated a more stable behavior due to the usage of multiscale kernels and in some cases followed the simulation data closer. Neither model was able to fully capture nonlinearities. We justified this result from the fact that the surrogates are attempting prediction of a boundary quantity that is dependent on the combustion process, but none of their far field inputs they are trained on fully carry this multiscale information. The fact that both of these models are able to achieve an error $<20\%$ when predicting multiscale boundary quantities just from far field inputs is significant, particulary considering that they take a few $ms$ to run compared to 24 hours for the DNS.
Second, we used the HMS in a forward UQ process to propagate uncertainty from the inputs to the QoI which is the regression rate. Assuming default priors for the input parameters (normal or uniform distributions as appropriate), we found that in the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.022$ mm/s, in the middle of the slab $\mu_{m}=0.081$ mm/s, and in the rear $\mu_{b}=0.147$ mm/s. The front and rear distributions appear to have a more central tendency compared to the middle part where the density is more evenly distributed on a wider range of values. We also found that the surrogates (and by extension the 2D DNS) do not correctly predict the high regression rate in the front ramp of the slab burner.
Last, we used the Bayesian approach to calibrate two of the parameter inputs, the latent heat of sublimation $l_{v}$ and the temperature exponent $b_{249}$ with available experimental data. We found the posterior distribution of $l_{v}$ to have statistics $\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$, which correspond to values that are much lower than the default value. The resulting posterior distribution for $b_{249}$ ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) includes similar values compared to the default value 2.54 from the DNS. We also conducted a study to compare the results of forward UQ when using the previous priors for the parameters and calibrated priors, for a fixed inlet velocity corresponding to the experiment conditions. We found that the two calibrated parameters have to compensate for the model inadequacy between the 2D DNS and the experimental data and their effect is not sufficient to make the surrogate be accurate in all slab locations simultaneously. When uncalibrated, the surrogate underestimates the regression rate in the front part, correctly predicts the middle, and overestimates the rear part of the slab. When calibrated priors are used for the parameters, the surrogate correctly predicts the front, but further overestimates the middle and rear part of the slab. Overall, the conclusion is that there is not one set of ($l_{v},b_{249}$) that result in correctly predicting the regression rate in all locations of the slab.
In the future we aim to expand on the work shown here through various efforts. Considering the combustion of higher alkanes ($C_{n}H_{2n+2}$) such as paraffin wax in the same slab burner setup can lead to better understanding of hybrid rocket systems. Fuels like paraffin are also of interest due to higher regression rate compared to other polymers, and therefore UQ of the regression rate from the DNS and validation with equivalent experiments are also relevant. The need to investigate surrogates that are able to produce accurate estimates of time-varying field QoIs is also of interest, specifically when these surrogates need to be coupled as parts of sequential frameworks. Thus, future work can explore applications of our framework where a correlated, joint PDF of parameters estimated by fitting experimental or first-principles density-functional theory (DFT) calculations data using Bayesian inference techniques [72, 88] [89]. Lastly, a more in depth calibration effort with more experimental data would help understand better how the surrogates perform across a wide range of conditions as well as identify the conditions that surrogates are appropriate for.
Acknowledgements
This work was supported by the United States Department of Energy’s (DoE) National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program III (PSAAP III) at the University at Buffalo, under contract number DE-NA0003961.
The authors would also like to acknowledge the Tufts University High Performance Compute Cluster which was used for running the uncertainty analyses.
We thank Kolos Retfalvi, Jasper Stedman, Alex Post, and Joseph Cygan from the University at Buffalo for their help with acquiring and tracing the experimental data used for the Bayesian calibration.